# Xing Huang a), Feng-Yu Wang a)

### v2

@@ 3. $(A_{2})$
(2.29) $\begin{split}&\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon+r},P_{t}^{*}\gamma^{\varepsilon})^{2}=\sup_{[f]_{\alpha}\leq 1}\big{|}P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})\big{|}^{2}\leq I_{1}+I_{2},\\ &I_{1}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{|}^{2},\\ &I_{2}:=2\bigg{|}\mathbb{E}\int_{0}^{r}\frac{\text{\rm{d}}}{\text{\rm{d}}\theta}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\text{\rm{d}}\theta\bigg{|}^{2}=2\bigg{|}\mathbb{E}\int_{0}^{r}\Big{\{}\nabla_{X_{0}^{\tilde{\gamma}}-X_{0}^{\gamma}}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\Big{\}}\text{\rm{d}}\theta\bigg{|}^{2}.\end{split}$
@@
(2.30) $\begin{split}&\mathbb{E}|R_{t}^{\varepsilon,r}-1|^{2}=\mathbb{E}\big{[}(R_{t}^{\varepsilon,r})^{2}-1\big{]}\leq\mathrm{esssup}_{\Omega}(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}-1)\\ &\leq\mathrm{esssup}_{\Omega}\left(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s\right)\\ &\leq\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,\end{split}$
@@
Combining this with (2.1), (2.15), (2.30) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, where $\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$ is defined in (2.12) with $\gamma^{\varepsilon}$ replacing $\gamma$, we can find constants $k_{1},k_{2}>1$ such that
@@
Combining this with Lemma 2.3(1), (2.15), (A) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, we find constants $c_{5}>0$ such that
@@
(2.34) $\begin{split}I_{2}&\leq 2\bigg{(}\mathbb{E}\bigg{[}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|\int_{0}^{r}|\nabla P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})|\text{\rm{d}}\theta\bigg{]}\bigg{)}^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t}r^{2}(\mathbb{E}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|)^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}r^{2}}{t}\big{(}\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}]\big{)}^{\frac{2}{k}}.\end{split}$
@@
Noting that $\gamma^{1}=\tilde{\gamma}$, this implies the desired estimate. ∎
@@
Let $k=2$. According to [28, Theorem 2.5] for $D=\mathbb{R}^{d}$, see also [32, Theorem 4.1], (A) implies the following log-Harnack inequality for some constant $c_{0}>0$ and any $\gamma\in\mathscr{P}_{2}$:
@@
Then by [28, (4.13)], see also [13, Theorem 2.1], it suffices to find a constant $c>0$ such that
@@ 3 Bismut Formula
Let $k\in(1,\infty)$ and denote $k^{*}:=\frac{k}{k-1}$. In this part, we establish the Bismut formula for the intrinsic derivative of $P_{t}f(\gamma)$ for $\gamma\in\mathscr{P}_{k}.$ To this end, we assume
(3.1) $b_{t}(x,\mu)=b_{t}^{(0)}(x)+B_{t}(x,\mu,\mu(V)),\ \ t\in[0,T],x\in\mathbb{R}^{d},\mu\in\mathscr{P},$
where for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$,
$V:\mathbb{R}^{d}\rightarrow\mathbb{B},\ \ \ \ B:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\times\mathbb{B}\rightarrow\mathbb{R}^{d}$
are measurable such that $[V]_{\alpha}\leq 1$ for some $\alpha\in\mathscr{A}$, i.e. $V$ is only square root Dini continuous and hence $B_{t}(x,\mu,\mu(V))$ may be not intrinsically differentiable in $\mu$, which produces essential difficulty. The following definition is taken from [3] using the idea of [1].
@@
If moreover
$\lim_{\|\phi\|_{T_{\mu,k}}\downarrow 0}\frac{\|f(\mu\circ(id+\phi)^{-1})-f(\mu)-D_{\phi}^{I}f(\mu)\|_{\mathbb{B}}}{\|\phi\|_{T_{\mu,k}}}=0,$
then $f$ is called $L$-differentiable at $\mu$. The function $f$ is called intrinsically differentiable or $L$-differentiable on $\mathscr{P}_{k}$, if it is so at any $\mu\in\mathscr{P}_{k}$.
Recall that a real function $f$ on a Banach space $\mathbb{B}$ is called Gateaux differentiable, if for any $z\in\mathbb{B}$,
$v\mapsto\nabla_{v}^{\mathbb{B}}f(z):=\lim_{\varepsilon\downarrow 0}\frac{f(z+\varepsilon v)-f(z)}{\varepsilon}$
is a well-defined bounded linear functional. In this case we denote
$\|\nabla^{\mathbb{B}}f(z)\|_{\mathbb{B}^{*}}:=\sup_{\|v\|_{\mathbb{B}}\leq 1}|\nabla_{v}^{\mathbb{B}}f(z)|.$
Moreover, $f$ is called Fréchet differentiable if it is Gateaux differentiable and
$\lim_{\|v\|_{\mathbb{B}}\downarrow 0}\frac{|f(z+v)-f(z)-\nabla_{v}^{\mathbb{B}}f(z)|}{\|v\|_{\mathbb{B}}}=0,\ \ z\in\mathbb{B}.$
It is well-known that a Gateaux differentiable function $f$ is Fréchet differentiable provided $\nabla_{v}^{\mathbb{B}}f(z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$.
When $\mathbb{B}=\mathbb{R}^{l}$ for some $l\geq 1$, we simply denote $\nabla^{\mathbb{B}}=\nabla$.
@@ 3. $(B_{2})$
    There exists $\alpha\in\mathscr{A}$ and $K_{V}\geq 0$ such that
    $[V]_{\alpha}:=\sup_{x\neq y}\frac{\|V(x)-V(y)\|_{\mathbb{B}}}{\alpha(|x-y|)}\leq 1,\ \ \|V(x)\|_{\mathbb{B}}\leq K_{V}(1+|x|^{k-1}),\ \ x\in\mathbb{R}^{d}.$
@@ 4. $(B_{3})$
    For any $t\in[0,T]$, we have $B_{t}\in C(\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B})$. $B_{t}(x,\mu,z)$ is differentiable in $x$, $L$-differentiable in $\mu\in\mathscr{P}_{k}$, and Fréchet differentiable in $z\in\mathbb{B}$ such that $\nabla_{v}^{\mathbb{B}}B_{t}(x,\mu,z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$. Next, there exists a constant $K>0$ such that
    $|B_{t}(0,\delta_{0},0)|\leq K,$
    $|\nabla B_{t}(\cdot,\mu,z)(x)|+\|D^{L}B_{t}(x,\cdot,z)(\mu)\|_{L^{k^{*}}(\mu)}+\|\nabla^{\mathbb{B}}B_{t}(x,\mu,\cdot)(z)\|_{\mathbb{B}^{*}}\leq K,$
    $\ \ (t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}.$
    Moreover, for any $(t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}$, there exists a constant $c(t,x,\mu,z)>0$ such that
    |$|\{D^{L}B_{t}(x,\cdot,z)(\mu)(y)|\leq c(t,x,\mu,z)(1+|y|^{k-1}),\ \ y\in\mathbb{R}^{d}.$
@@
Let $l$ be a positive integer and let $\alpha(s)=s^{\varepsilon}$ for some $\varepsilon\in(0,1\wedge(k-1))$. Then $V:\mathbb{R}^{d}\rightarrow\mathbb{R}^{l}$ with
$|V(x)-V(y)|\leq\alpha(|x-y|),\ \ x,y\in\mathbb{R}^{d}$
satisfies ($B_{2}$). Moreover, let $g\in C^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{l};\mathbb{R}^{d})$ with bounded gradient. Then $B_{t}(x,\mu,z)=\int_{\mathbb{R}^{d}}g(x,y,z)\mu(\text{\rm{d}}y)$ satisfies ($B_{3}$).
@@
(3.2) $\begin{split}&\text{\rm{d}}X_{t}^{x,\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\mu})+B_{t}(X_{t}^{x,\mu},P_{t}^{*}\mu,P_{t}V(\mu))\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\mu})\text{\rm{d}}W_{t},\\ &\qquad\qquad\qquad X_{0}^{x,\mu}=x,\ \ t\in[0,T].\end{split}$

### v4

@@ 2. $(2)$
(2.34) $\begin{split}&\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon+r},P_{t}^{*}\gamma^{\varepsilon})^{2}=\sup_{[f]_{\alpha}\leq 1}\big{|}P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})\big{|}^{2}\leq I_{1}+I_{2},\\ &I_{1}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{|}^{2},\\ &I_{2}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\mathbb{E}\big{[}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+r}})-P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon}})\big{]}\bigg{|}^{2}.\end{split}$
@@
$\mathbb{E}|R_{t}^{\varepsilon,r}-1|^{2}=\mathbb{E}\big{[}(R_{t}^{\varepsilon,r})^{2}-1\big{]}\leq\mathrm{esssup}_{\Omega}(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}-1)$
(2.35) $\leq\mathrm{esssup}_{\Omega}\left(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s\right)$
$\leq\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,$
@@
Combining this with (2.1), (2.18), (2.2) and (2.10) with $z=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, where $\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$ is defined in (2.16) with $\gamma^{\varepsilon}$ replacing $\gamma$, we can find constants $k_{1},k_{2}>1$ such that
@@
By (2.19), we find a constant $c_{5}>0$ such that
@@
$I_{2}$ $\leq 2\sup_{[f]_{\alpha}\leq 1}\bigg{(}\mathbb{E}\bigg{[}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|\int_{0}^{r}|\nabla P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})|\text{\rm{d}}\theta\bigg{]}\bigg{)}^{2}$
(2.39) $\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t}r^{2}(\mathbb{E}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|)^{2}$
$\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}r^{2}}{t}\big{(}\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}]\big{)}^{\frac{2}{k}}.$
@@
This implies (2.28) by noting that $\gamma^{1}=\tilde{\gamma}$ and $\sup_{t\in[0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\leq\frac{\sqrt{T}\vee 1}{\alpha(1)}$ due to (2.2) . ∎
@@
Let $k=2$. According to [29, Theorem 2.5] for $D=\mathbb{R}^{d}$, see also [33, Theorem 4.1], (A) implies the following log-Harnack inequality for some constant $c_{0}>0$ and any $\gamma\in\mathscr{P}_{2}$:
@@
Then by [29, (4.13)], see also [13, Theorem 2.1], it suffices to find a constant $c>0$ such that
@@ 3 Bismut Formula
Let $k\in(1,\infty)$ and denote $k^{*}:=\frac{k}{k-1}$. In this part, we consider the SDE (1.1) with $(\tilde{\mathscr{P}},\tilde{\rho})=(\mathscr{P}_{k},\mathbb{W}_{k})$, where as in $(A_{2})$ the drift $b$ is decomposed as
(3.1) $b_{t}(x,\nu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\nu),\ \ t\in[0,T],x\in\mathbb{R}^{d},\ \nu\in\mathscr{P}_{k}.$
We aim to establish Bismut type formula for the intrinsic derivative of $\mathscr{P}_{k}\ni\mu\mapsto P_{t}f(\mu)$ for bounded measurable functions $f$ on $\mathbb{R}^{d}$, by only assuming that the extrinsic derivative $D^{E}b_{t}(x,\mu)(z)$ of the drift has a half-Dini continuity in $z\in\mathbb{R}^{d}$.
To this end, we first recall the notions of intrinsic and extrinsic derivatives which go back to [1], see [3] and [21].
@@
The function $f$ is called intrinsically differentiable on $\mathscr{P}_{k}$, if it is so at any $\mu\in\mathscr{P}_{k}$.
Next, we recall the (convexity) extrinsic derivative, see e.g. [21, Definition 1.2].
Definition 3.2.
A real function $f$ on $\mathscr{P}_{k}$ is called extrinsically differentiable on $\mathscr{P}_{k}$ with derivative $D^{E}f$ if
$D^{E}f(\mu)(x)=\lim_{\varepsilon\downarrow 0}\frac{f((1-\varepsilon)\mu+\varepsilon\delta_{x})-f(\mu)}{\varepsilon}\in\mathbb{R}$
exists for all $(x,\mu)\in\mathbb{R}^{d}\times\mathscr{P}_{k}$. When $f=(f^{1},f^{2},\cdots,f^{d})$ is an $\mathbb{R}^{d}$-valued function on $\mathscr{P}_{k}$, we denote $D^{E}f=(D^{E}f^{1},D^{E}f^{2},\cdots,D^{E}f^{d})$.
@@ 3. $(B_{2})$
    For any $t\in[0,T],y\in\mathbb{R}^{d}$, $b_{t}^{(1)}(y,\cdot)$ is extrinsically differentiable in $\mathscr{P}_{k}$ with the extrinsic derivative $D^{E}b_{t}^{(1)}(y,\nu)(z)$ being continuous in $(y,\nu,z)\in\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{R}^{d}$. Moreover, there exists $\alpha\in\mathscr{A}$ with $\alpha\leq c_{0}(1+|\cdot|^{k-1})$ for some $c_{0}>0$ such that
    $|D^{E}b_{t}^{(1)}(y,\nu)(z)-D^{E}b_{t}^{(1)}(y,\nu)(\bar{z})|\leq\alpha(|z-\bar{z}|),$
    $\qquad\quad z,\bar{z}\in\mathbb{R}^{d},t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
@@ 4. $(B_{3})$
    For any $t\in[0,T]$, $\nu\in\mathscr{P}_{k}$, $b_{t}^{(1)}(\cdot,\nu)$ is differentiable and there exists a constant $\tilde{K}>0$ such that
    $|b_{t}^{(1)}(0,\delta_{0})|\leq\tilde{K},\ \ |\nabla b_{t}^{(1)}(y,\nu)|\leq\tilde{K},\ \ (t,y,\nu)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}.$
As indicated in Introduction that existing results on Bismut type formulas for the intrinsic derivative of $P_{t}f(\mu)$ are established under upper bound conditions on the $L$-derivative of $b_{t}(y,\nu)$ in $\nu$. Noting that under a mild condition, the $L$-derivative equals to the gradient of the extrinsic derivative, so the above condition on the $\alpha$-continuity of $D^{E}b_{t}^{(1)}(y,\nu)(z)$ in $z$ is much weaker. To see this, we present below a simple example.
@@
Let $\alpha(s)=s^{\varepsilon}$ for some $\varepsilon\in(0,1\wedge(k-1))$. Let $g:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfy
$|g(y,z)-g(y,\bar{z})|\leq\alpha(|z-\bar{z}|),\ \ |\nabla g(\cdot,z)|\leq K,\ \ y,z,\bar{z}\in\mathbb{R}^{d}$
for some $K\geq 0$. Let $b_{t}^{(1)}(y,\nu)=\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z)$. By Definition 3.2, it holds
$D^{E}b_{t}^{(1)}(y,\nu)(z)=g(y,z)-\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z),\ \ y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k},z\in\mathbb{R}^{d}.$
However, by Definition 3.1, $b_{t}^{(1)}(y,\nu)$ is not intrinsically differentiable in $\nu$. In fact, since $g(y,z)$ is not differentiable in $z$, for any $y\in\mathbb{R}^{d}$, $\nu\in\mathscr{P}_{k}$, $\phi\in L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\nu)$, the limit
$\lim_{r\downarrow 0}\frac{\int_{\mathbb{R}^{d}}g(y,z+r\phi(z))\nu(\text{\rm{d}}z)-\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z)}{r}$
does not exist. Moreover, it holds
$|D^{E}b_{t}^{(1)}(y,\nu)(z)-D^{E}b_{t}^{(1)}(\bar{y},\bar{\nu})(\bar{z})|$
$=|g(y,z)-g(\bar{y},\bar{z})|+\left|\int_{\mathbb{R}^{d}}g(y,z)\nu(\text{\rm{d}}z)-\int_{\mathbb{R}^{d}}g(\bar{y},z)\bar{\nu}(\text{\rm{d}}z)\right|$
$\leq\alpha(|z-\bar{z}|)+2K|y-\bar{y}|+\mathbb{W}_{\alpha}(\nu,\bar{\nu}),\ \ y,\bar{y}\in\mathbb{R}^{d},\nu,\bar{\nu}\in\mathscr{P}_{k},z,\bar{z}\in\mathbb{R}^{d}.$
Note that Jensen’s inequality implies that
$\mathbb{W}_{\alpha}(\mu,\nu)\leq\inf_{\pi\in\mathscr{C}(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\alpha(|x-y|)\pi(\text{\rm{d}}x,\text{\rm{d}}y)\leq\alpha(\mathbb{W}_{1}(\mu,\nu))\leq\alpha(\mathbb{W}_{k}(\mu,\nu)),\ \ \mu,\nu\in\mathscr{P}_{k}.$
So, $D^{E}b_{t}^{(1)}(y,\nu)(z)$ is continuous in $(y,\nu,z)\in\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{R}^{d}$. Finally, by the dominated convergence theorem, we have
$\nabla b_{t}^{(1)}(\cdot,\nu)=\int_{\mathbb{R}^{d}}\nabla g(\cdot,z)\nu(\text{\rm{d}}z),\ \ \nu\in\mathscr{P}_{k}.$
Therefore, $b^{(1)}$ satisfies $(B_{2})$-$(B_{3})$.
@@
(3.2) $\begin{split}&\text{\rm{d}}X_{t}^{x,\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\mu})+b_{t}^{(1)}(X_{t}^{x,\mu},P_{t}^{*}\mu)\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\mu})\text{\rm{d}}W_{t},\\ &\qquad\qquad\qquad X_{0}^{x,\mu}=x,\ \ t\in[0,T].\end{split}$
Let
$\mathscr{B}_{k,b}(\mathbb{R}^{d}):=\bigg{\{}f:\ \frac{f}{1+|\cdot|^{k}}\in\mathscr{B}_{b}(\mathbb{R}^{d})\bigg{\}}.$
We first give a lemma on Bismut formula of $P_{t}^{\mu}f$ for $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$.
Lemma 3.2.
Let $\sigma$ and $b$ satisfy (B). Then for any $v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, the limit
Moreover, the Bismut formula for $P^{\gamma}_{t}$ holds:
(3.4) $\begin{split}&\nabla_{v}P^{\gamma}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right],\\ &\qquad\zeta_{s}:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1},\ \ \ f\in\mathscr{B}_{k,b}(\mathbb{R}^{d}),\ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].\end{split}$
Proof.
By [30, Theorem 2.1] for $\beta_{s}=\frac{s}{t}$, (B) implies (3.3) and (3.4) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$. To deduce (3.4) for any $f\in\mathscr{B}_{b,k}(\mathbb{R}^{d}),$ let
$f_{n}:=[(-n)\lor f]\land n,\ \ n\geq 1.$
By (2.8), (3.3) and the boundedness of $\zeta_{s}$, we find constants $c_{0},c_{1}(\gamma)>0$ such that
(3.5) $\begin{split}&\mathbb{E}\bigg{[}\big{(}1+|X_{t}^{x+rv,\gamma}|^{k}\big{)}\bigg{|}\int_{0}^{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\bigg{|}\bigg{]}\\ &\leq c_{0}\sqrt{t}\Big{(}\mathbb{E}\big{[}2+|X_{t}^{x+rv,\gamma}|^{2k}\big{]}\Big{)}^{\frac{1}{2}}\Big{(}\mathbb{E}\big{[}\sup_{s\in[0,T]}\big{|}\nabla_{v}X^{x+rv,\gamma}_{s}\big{|}^{2}\big{]}\Big{)}^{\frac{1}{2}}\\ &\leq c_{1}(\gamma)\sqrt{t}|v|\big{(}1+|x|^{k}+|v|^{k}\big{)},\ \ t\in(0,T],x,v\in\mathbb{R}^{d},r\in[0,1].\end{split}$
By (3.4) for $f_{n}$ in place of $f$, we obtain
$\frac{P^{\gamma}_{t}f_{n}(x+\varepsilon v)-P^{\gamma}_{t}f_{n}(x)}{\varepsilon}$
$=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\mathbb{E}\left[f_{n}(X^{x+rv,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\text{\rm{d}}r.$
Since $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$, by (2.8), (3.3) and (3.5), we may apply the dominated convergence theorem such that the above formula with $n\rightarrow\infty$ implies
(3.6) $\begin{split}&\frac{P^{\gamma}_{t}f(x+\varepsilon v)-P^{\gamma}_{t}f(x)}{\varepsilon}\\ &=\frac{1}{\varepsilon}\int_{0}^{\varepsilon}\mathbb{E}\left[f(X^{x+rv,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\text{\rm{d}}r,\\ &\qquad f\in\mathscr{B}_{k,b}(\mathbb{R}^{d}),\varepsilon>0,\ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].\end{split}$
Note that (3.4) for $f_{n}$ in place of $f$ yields
$\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}f_{n}(x+\varepsilon v)-P^{\gamma}_{t}f_{n}(x)}{\varepsilon}-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}$
$=\lim_{n\rightarrow\infty}\bigg{|}\mathbb{E}\left[f_{n}(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]$
$\qquad\qquad\quad-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}=0,$
where the last step follows from the dominated convergence theorem due to $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$ and (3.5). This together with (3.6) for $f-f_{n}$ in place of $f$ implies
(3.7) $\begin{split}&\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}f(x+\varepsilon v)-P^{\gamma}_{t}f(x)}{\varepsilon}-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}\\ &\leq\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}(f-f_{n})(x+\varepsilon v)-P^{\gamma}_{t}(f-f_{n})(x)}{\varepsilon}\bigg{|}\\ &+\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\bigg{|}\frac{P^{\gamma}_{t}f_{n}(x+\varepsilon v)-P^{\gamma}_{t}f_{n}(x)}{\varepsilon}\\ &\qquad\qquad\qquad\qquad\quad-\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right]\bigg{|}\\ &\leq\lim_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\frac{1}{t\varepsilon}\int_{0}^{\varepsilon}\mathbb{E}\bigg{|}(f_{n}-f)(X_{t}^{x+rv,\gamma})\\ &\qquad\qquad\qquad\qquad\qquad\quad\times\int_{0}^{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\bigg{|}\text{\rm{d}}r.\end{split}$
Since $f\in\mathscr{B}_{k,b}(\mathbb{R}^{d})$ implies
$|(f_{n}-f)(x)|\leq c(1+|x|^{k})1_{\{c(1+|x|^{k})\geq n\}},\ \ n\geq 1$
for some constant $c>0$, by the same reason leading to (3.5), we find constants $\tilde{c},c_{2}(\gamma)>0$ such that
$\sup_{r\in[0,1]}\mathbb{E}\bigg{|}(f_{n}-f)(X_{t}^{x+rv,\gamma})\int_{0}^{t}\big{\langle}\zeta_{s}(X^{x+rv,\gamma}_{s})\nabla_{v}X^{x+rv,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\bigg{|}$
$\leq\tilde{c}\sqrt{t}|v|\sup_{r\in[0,1]}\Big{(}\mathbb{E}\big{[}1+|X_{t}^{x+rv,\gamma}|^{4k}\big{]}\Big{)}^{\frac{1}{2}}n^{-1}$
$\leq c_{2}(\gamma)\sqrt{t}|v|\big{(}1+|x|^{2k}+|v|^{2k}\big{)}n^{-1}.$
Therefore, (3.4) follows form (3.7). ∎